# Maël Dumas LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Anthony Perez LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Mathis Rocton Algorithms and Complexity Group, TU Wien, Vienna, Austria Ioan Todinca LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France

### v3

@@ 2.1 Hardness results
We first show the NP-completeness of Block Graph Deletion and Block Graph Editing by giving a reduction from Cluster Deletion and Cluster Editing, known to be NP-complete [34, 39, 36]. Notice that Block Graph Completion admits a polynomial-time algorithm due to the very nature of block graphs: any biconnected component of the input graph (computable in linear time [28]) must be turned into a clique.
Lemma 6.
@@
First, suppose that $H^{\prime}=G^{\prime}-F^{\prime}$ is not connected. Let $C$ be the connected component that contains $u$. For any other connected component $C^{\prime}$ of $H^{\prime}$, take a vertex $v\in V(C^{\prime})$ and construct the deletion set $F^{\prime\prime}=F^{\prime}\setminus\{(u,v)\}$. By Lemma 3, $(C,\{u\})\otimes(C^{\prime},\{v\})$ is a block graph, and the other connected components of $H^{\prime}$ are unchanged. Hence $G^{\prime}-F^{\prime\prime}$ is a block graph and since $|F^{\prime\prime}|<|F^{\prime}|$, $F^{\prime\prime}$ is a $k$-deletion. We iterate this construction until we get an intermediate solution $F^{*}_{1}$ such that $H^{*}_{1}=G^{\prime}-F^{*}_{1}$ is connected.
If $F^{*}_{1}$ does not affect $u$ we set $F^{*}=F^{*}_{1}$ and we are done. Otherwise there is at least one vertex $v$ not adjacent to $u$ in $H^{*}_{1}$. Since $H^{*}_{1}$ is connected we can choose $v$ at distance exactly $2$ from $u$. Since $H^{*}_{1}$ is a block graph we have $N(u)\cap N(v)=\{w\}$. Consider the biconnected component $C$ containing $v$ and $w$, and let $C^{\prime}=C\setminus\{w\}$. Since $H^{*}_{1}$ is a block graph, $C$ (and thus $C^{\prime}$) is a clique. Let $K_{u}$ and $K_{v}$ be the connected components of $H^{*}_{1}-(\{w\}\times C^{\prime})$ that contain the vertices $u$ and $v$ respectively, they are block graphs by heredity. By Lemma 3 the graph $H^{*}_{2}=(K_{u},\{u\})\otimes(K_{v},\{C^{\prime}\})$ is a block graph. Observe that ($(\{u\}\times C^{\prime})\subseteq F^{*}_{1}$ since $u$ is an universal vertex and let $F^{*}_{2}$ be the deletion set such that $H^{*}_{2}=G^{\prime}-F^{*}_{2}$. By construction we can observe that $|F^{*}_{2}|=|F^{*}_{1}|$, hence $F^{*}_{2}$ is a $k$-deletion of $G^{\prime}$ into a block graph. We iterate this construction until we get a $k$-deletion $F^{*}$ of $G$ that does not affect $u$.
@@
The same reduction can be done for Block Graph Editing, one needs only to observe that there are no added edges incident to $u$, and we can use the same argument as above on the deleted edges: if there is no deleted edge incident to $u$ we are done, otherwise we construct another set of deleted edges of equal size, but containing fewer edges incident to $u$. If in the construction we have to delete an edge that was added by the edition, it is clear that removing this edge from the edition set result in a strictly smaller edition set. ∎
The NP-completeness of Strictly Chordal Completion follows directly from the proof of NP-completeness of 3-Leaf Power Completion from [16, Theorem 3].
Lemma 7.
@@
The same reduction can be done from Cluster Deletion to Strictly Chordal Deletion. This concludes the proof. ∎
@@ 3 Kernelization algorithm for Block Graph Editing
We begin this section by providing a high-level description of our kernelization algorithm. A similar technique will be applied for modification problems towards strictly chordal graphs but on the critical clique graph rather than the original one. Let us consider a positive instance $(G=(V,E),k)$ of Block Graph Editing, $F$ a suitable solution and $H=G\triangle F$. Since $|F|\leqslant k$, we know that at most $2k$ vertices of $H$ may be affected vertices. Let $A$ be the set of such vertices, $T$ the minimum induced subgraph of $H$ that spans all vertices of $A$ and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$. From Lemma 5 we have $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$. We will define the notion of BG-branch, corresponding to subgraphs of $G$ that induce block graphs. We will focus our interest on two types of BG-branches: the ones that are connected to the rest of the graph by only one (cut) vertex, called $1$-BG-branches, and the ones that are connected to the rest of the graph by exactly two non-adjacent vertices, called $2$-BG-branches. We will prove that we can keep only the cut-vertices of $1$-BG-branches and reduce the $2$-BG-branches containing more than $k+3$ vertices. We will see that there are two kinds of connected components remaining in the graph $H\setminus(A\cup A^{\prime})$, the ones adjacent to maximal cliques of vertices of $A\cup A^{\prime}$, corresponding to critical cliques (which can be bounded to a linear number of vertices) and the ones adjacent to two non-adjacent vertices of $A\cup A^{\prime}$ (which correspond to $2$-BG-branches in $H$). Since $H$ is a block graph, we know that there is at most $|A\cup A^{\prime}|=O(k)$ maximal cliques of vertices of $A\cup A^{\prime}$. From Lemma 5 we have that there are at most $4k$ connected components in $T\setminus(A\cup A^{\prime})$, and each one corresponds to a $2$-BG-branch in the graph $H$. It remains that $H\setminus(A\cup A^{\prime})$ contains a linear number of connected components, each one containing $O(k)$ vertices. Altogether, the graph $H$ contains $O(k^{2})$ vertices.
@@ 3.1 Classical reduction rules
We now turn our attention to Rule 3.2. Let $G^{\prime}$ denote the graph obtained from the removal of $|K|-(2k+2)$ arbitrary vertices from $K$ in $G$ and $K^{\prime}=V(G)\setminus V(G^{\prime})$. Assume first that $G$ admits a $k$-edition $F$. Observe that $F$ is a $k$-edition of $G^{\prime}$ by heredity. Conversely, assume that $G^{\prime}$ admits a $k$-edition $F$ and let $H^{\prime}=G^{\prime}\triangle F$. Notice that since $F$ affects at most $2k$ vertices and since $|K\setminus K^{\prime}|=2k+2$, there exist two unaffected vertices $u$ and $v$ in $K\setminus K^{\prime}$. We claim that $H=G\triangle F^{\prime}$ is a block graph. Assume for a contradiction that $H$ contains an obstruction $W$. By construction, $W$ must intersect $K^{\prime}$ since otherwise $W$ would exist in $H^{\prime}$. Moreover, notice that $|W\cap K^{\prime}|\leqslant 2$, otherwise $W$ would contain a clique module with $3$ or $4$ vertices, which is impossible in a cycle or a diamond. Hence assume first that $|W\cap K^{\prime}|=\{a\}$: in this case, since $u$ is unaffected we have $N_{H}[a]\setminus K^{\prime}=N_{H^{\prime}}[u]$ and thus the set $(W\setminus\{a\})\cup\{u\}$ induces an obstruction in $H^{\prime}$, a contradiction. Similarly, if $|W\cap K^{\prime}|=\{a,b\}$ the set $(W\setminus\{a,b\})\cup\{u,v\}$ induces an obstruction in $H^{\prime}$, leading once again to a contradiction. Finally, since the biconnected components can be computed in $O(n+m)$ time [28], since block graphs can be recognized in $O(n+m)$ time and since true twins can be detected in $O(n+m)$ time (using modular decomposition [38]), both rules can be computed in linear time. ∎
@@ 3.2 Reducing Block Graph branches
Conversely, let $F^{\prime}$ be a $k$-edition of $G^{\prime}$ into a block graph. Let $u$ be the attachment point of $B$ in $G$. Observe that since $B^{R}$ is connected and not empty, $u$ is not a cut-vertex of the block graph $B$ and its neighborhood in $B$ is either a maximal clique or a single vertex. Since $H^{\prime}=G^{\prime}\triangle F^{\prime}$ and $B$ are block graphs and $N_{B}(u)$ is a single vertex or a maximal clique of $B$, Lemma 3 ensures that $H=(H^{\prime},\{u\})\otimes(B^{R},N_{B}(u))$ is a block graph. We can observe that $H=G\triangle F^{\prime}$, hence $F^{\prime}$ is a $k$-edition of $G$ into a block graph. This concludes the proof. ∎
@@
Figure 4: Illustration of the case where $p_{1}$ and $p_{2}$ are not in the same connected component of $H^{\prime}$ in the Lemma 11. The white circles correspond to critical cliques of $B$. The blue sets correspond to $N_{B}(P_{1})$ and $N_{B}(P_{2})$.
Let $F$ be a $k$-edition of $G$ into a block graph, $H=G\triangle F$ and $H^{\prime}=H\setminus V(B^{R})$. First consider the case in which $p_{1}$ and $p_{2}$ are in two different connected components in $H^{\prime}$. Let $H^{\prime}_{1},H^{\prime}_{2}$ be the two connected components of $H^{\prime}$ containing $p_{1}$ and $p_{2}$, respectively, and $H^{\prime}_{3}$ the remaining of $H^{\prime}$. It follows from Lemma 3 that the graph $H^{*}$ resulting from the disjoint union of
@@
We can assume that if $p_{1}$ and $p_{2}$ are in the same component of $H^{\prime}$, $F$ contains an edge-cut of $B$ or there would be a path $\pi_{B}$ from $p_{1}$ to $p_{2}$ in $B$ that still exists in $B\triangle F$. Let $M$ be a min-cut of $B$ and consider $B_{1}$ (resp. $B_{2}$) the connected component containing $p_{1}$ (resp. $p_{2}$) in $B\triangle M$. Both $B^{R}_{1}=B_{1}\setminus\{p_{1}\}$ and $B^{R}_{2}=B_{2}\setminus\{p_{2}\}$ are induced subgraphs (one of them possibly empty) of $B$ and block graphs by heredity. By Lemma 3, the following graph is a block graph:
@@
Figure 5: Illustration of the case where $p_{1}$ and $p_{2}$ are in the same connected component of $H^{\prime}$ and $F$ contains a min-cut of $B$ in Lemma 21. The white circles correspond to critical cliques of $B$. The blue sets correspond to $N_{B}(P_{1})$ and $N_{B}(P_{2})$. The red dotted edges between $B_{1}$ and $B_{2}$ are edges of an edge-cut of $B$ deleted by the edition.
@@
In all cases, there exists a $k$-edition $F^{*}$ of $G$ into a block graph such that the set pairs of $F^{*}$ containing a vertex of $V(B^{R})$ is either empty or exactly a min-cut of $B$. This concludes the proof. ∎
Rule 3.4.
@@
Let $F$ be a $k$-edition of $G$ that satisfies Lemma 11 and $H=G\triangle F$. If $F$ contains a min-cut of $B$, observe that $mc(B^{\prime})=mc(B)\leqslant k$ and consider $B_{i}^{\prime},i\in\{1,2\}$ the connected component of $B^{\prime}\triangle F_{m}$ containing $p_{i}$, where $F_{m}$ is a min-cut of $B^{\prime}$. By Lemma 3 the graph
@@
is a block graph. Let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$. Since $mc(B^{\prime})=mc(B)\leqslant k$, by construction $|F^{\prime}|=|F|$ , therefore $F^{\prime}$ is a $k$-edition of $G^{\prime}$. If $F$ does not contain a min-cut of $B$, then no vertex of $B^{R}$ is affected and $p_{1},p_{2}$ are not in the same connected component of $H\setminus V(B^{R})$. Let $H_{1}$ and $H_{2}$ be the connected components of $H\setminus V(B^{R})$ that contain respectively $p_{1}$ and $p_{2}$ and $H_{3}$ the remaining components of $H\setminus V(B^{R})$. By Lemma 3 the graph $H^{\prime}$ corresponding to the disjoint union of
@@
and $H_{3}$ is a block graph. We can observe that $H^{\prime}=G^{\prime}\triangle F$, hence $F$ is a $k$-edition of $G^{\prime}$ into a block graph. The other way is symmetrical. ∎
@@ 3.3 Bounding the size of the kernel
Proof.
We show that we can enumerate all $1$-BG-branches and $2$-BG-branches in polynomial time. Recall that recognition of block graphs can be computed in linear time. All $1$-BG-branches can be enumerated by removing a vertex $v\in V(G)$ and looking among the remaining connected components those that induce a connected block graph together with $v$. We proceed similarly to detect clean $2$-BG-branches by removing a pair of non-adjacent vertices $u,v\in V(G)$ and looking among the remaining connected components those that induce a connected block graph together with $u$ and $v$. ∎
@@
Proof.
Let $(G=(V,E),k)$ be a reduced yes-instance of Block Graph Editing, $F$ a $k$-edition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. Let $A$ be the set of affected vertices of $H$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T_{A}$ be the connected minimal induced subgraph of $H$ that spans $A$, and $d(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$. By Lemma 5, $|d(T_{A})\setminus A|\leqslant 3\cdot|A|\leqslant 6k$.
@@
Now consider the connected components of type 2. We can observe that such a component $K$ together with its adjacent vertices $x,y\in A\cup d(T_{A})$ form a clean 2-$BG$-branch, and thus contains at most $32k+$ 3 vertices. By Lemma 5, $T_{A}\setminus(A\cup d(T_{A}))$ contains at most $2\cdot|A|=4k$ connected components which is exactly the number of type 2 components, hence there are at most $4k\cdot(2k+1)$ vertices in connected components of type 2.
We conclude that if $(G,k)$ is a reduced yes-instance, then $|V(G)|\leqslant 24k^{2}+28k$. Therefore, given a reduced instance $(G,k)$ of Block Graph Editing, if $|V(G)|>24k^{2}+28k$ we can safely conclude that $(G,k)$ is a no-instance and return a trivial no-instance of constant size, thus proving the size of our kernel. Finally we claim that the reduced instance can be computed in polynomial time. Indeed, all reduction rules can be exhaustively applied in polynomial time (Lemmata 9 and 13).
@@
For the case of Block Graph Deletion, all the lemmata in this section are still true. In particular Lemma 11 holds by simply replacing ”$k$-edition” with ”$k$-deletion”, and the proof for the deletion variant is contained in the edition one. Using the same set of rules, we have a kernel with $O(k^{2})$ vertices for Block Graph Deletion. ∎

### v4

@@ 2.1 Hardness results
We first show the NP-completeness of Block Graph Deletion and Block Graph Editing by giving a reduction from Cluster Deletion and Cluster Editing, known to be NP-complete (34, 39, 36). Notice that Block Graph Completion admits a polynomial-time algorithm due to the very nature of block graphs: any biconnected component of the input graph (computable in linear time (28)) must be turned into a clique.
Lemma 2.12.
@@
First, suppose that $H^{\prime}=G^{\prime}-F^{\prime}$ is not connected. Let $C$ be the connected component that contains $u$. For any other connected component $C^{\prime}$ of $H^{\prime}$, take a vertex $v\in V(C^{\prime})$ and construct the deletion set $F^{\prime\prime}=F^{\prime}\setminus\{(u,v)\}$. By Lemma 2.6, $(C,\{u\})\otimes(C^{\prime},\{v\})$ is a block graph, and the other connected components of $H^{\prime}$ are unchanged. Hence $G^{\prime}-F^{\prime\prime}$ is a block graph and since $|F^{\prime\prime}|<|F^{\prime}|$, $F^{\prime\prime}$ is a $k$-deletion. We iterate this construction until we get an intermediate solution $F^{*}_{1}$ such that $H^{*}_{1}=G^{\prime}-F^{*}_{1}$ is connected.
If $F^{*}_{1}$ does not affect $u$ we set $F^{*}=F^{*}_{1}$ and we are done. Otherwise there is at least one vertex $v$ not adjacent to $u$ in $H^{*}_{1}$. Since $H^{*}_{1}$ is connected we can choose $v$ at distance exactly $2$ from $u$. Since $H^{*}_{1}$ is a block graph we have $N(u)\cap N(v)=\{w\}$. Consider the biconnected component $C$ containing $v$ and $w$, and let $C^{\prime}=C\setminus\{w\}$. Since $H^{*}_{1}$ is a block graph, $C$ (and thus $C^{\prime}$) is a clique. Let $K_{u}$ and $K_{v}$ be the connected components of $H^{*}_{1}-(\{w\}\times C^{\prime})$ that contain the vertices $u$ and $v$ respectively, they are block graphs by heredity. By Lemma 2.6 the graph $H^{*}_{2}=(K_{u},\{u\})\otimes(K_{v},\{C^{\prime}\})$ is a block graph. Observe that ($(\{u\}\times C^{\prime})\subseteq F^{*}_{1}$ since $u$ is an universal vertex and let $F^{*}_{2}$ be the deletion set such that $H^{*}_{2}=G^{\prime}-F^{*}_{2}$. By construction we can observe that $|F^{*}_{2}|=|F^{*}_{1}|$, hence $F^{*}_{2}$ is a $k$-deletion of $G^{\prime}$ into a block graph. We iterate this construction until we get a $k$-deletion $F^{*}$ of $G$ that does not affect $u$.
@@
The same reduction can be done for Block Graph Editing, one needs only to observe that there are no added edges incident to $u$, and we can use the same argument as above on the deleted edges: if there is no deleted edge incident to $u$ we are done, otherwise we construct another set of deleted edges of equal size, but containing fewer edges incident to $u$. If in the construction we have to delete an edge that was added by the edition, it is clear that removing this edge from the edition set result in a strictly smaller edition set.
The NP-completeness of Strictly Chordal Completion follows directly from the proof of NP-completeness of 3-Leaf Power Completion from (16, Theorem 3).
Lemma 2.14.
@@
The same reduction can be done from Cluster Deletion to Strictly Chordal Deletion. This concludes the proof.
@@ 3 Kernelization algorithm for Block Graph Editing
We begin this section by providing a high-level description of our kernelization algorithm. A similar technique will be applied for modification problems towards strictly chordal graphs but on the critical clique graph rather than the original one. Let us consider a positive instance $(G=(V,E),k)$ of Block Graph Editing, $F$ a suitable solution and $H=G\triangle F$. Since $|F|\leqslant k$, we know that at most $2k$ vertices of $H$ may be affected vertices. Let $A$ be the set of such vertices, $T$ the minimum induced subgraph of $H$ that spans all vertices of $A$ and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$. From Lemma 2.10 we have $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$. We will define the notion of BG-branch, corresponding to subgraphs of $G$ that induce block graphs. We will focus our interest on two types of BG-branches: the ones that are connected to the rest of the graph by only one (cut) vertex, called $1$-BG-branches, and the ones that are connected to the rest of the graph by exactly two non-adjacent vertices, called $2$-BG-branches. We will prove that we can keep only the cut-vertices of $1$-BG-branches and reduce the $2$-BG-branches containing more than $k+3$ vertices. We will see that there are two kinds of connected components remaining in the graph $H\setminus(A\cup A^{\prime})$, the ones adjacent to maximal cliques of vertices of $A\cup A^{\prime}$, corresponding to critical cliques (which can be bounded to a linear number of vertices) and the ones adjacent to two non-adjacent vertices of $A\cup A^{\prime}$ (which correspond to $2$-BG-branches in $H$). Since $H$ is a block graph, we know that there is at most $|A\cup A^{\prime}|=O(k)$ maximal cliques of vertices of $A\cup A^{\prime}$. From Lemma 2.10 we have that there are at most $4k$ connected components in $T\setminus(A\cup A^{\prime})$, and each one corresponds to a $2$-BG-branch in the graph $H$. It remains that $H\setminus(A\cup A^{\prime})$ contains a linear number of connected components, each one containing $O(k)$ vertices. Altogether, the graph $H$ contains $O(k^{2})$ vertices.
@@ 3.1 Classical reduction rules
We now turn our attention to Rule 3.2. Let $G^{\prime}$ denote the graph obtained from the removal of $|K|-(2k+2)$ arbitrary vertices from $K$ in $G$ and $K^{\prime}=V(G)\setminus V(G^{\prime})$. Assume first that $G$ admits a $k$-edition $F$. Observe that $F$ is a $k$-edition of $G^{\prime}$ by heredity. Conversely, assume that $G^{\prime}$ admits a $k$-edition $F$ and let $H^{\prime}=G^{\prime}\triangle F$. Notice that since $F$ affects at most $2k$ vertices and since $|K\setminus K^{\prime}|=2k+2$, there exist two unaffected vertices $u$ and $v$ in $K\setminus K^{\prime}$. We claim that $H=G\triangle F^{\prime}$ is a block graph. Assume for a contradiction that $H$ contains an obstruction $W$. By construction, $W$ must intersect $K^{\prime}$ since otherwise $W$ would exist in $H^{\prime}$. Moreover, notice that $|W\cap K^{\prime}|\leqslant 2$, otherwise $W$ would contain a clique module with $3$ or $4$ vertices, which is impossible in a cycle or a diamond. Hence assume first that $|W\cap K^{\prime}|=\{a\}$: in this case, since $u$ is unaffected we have $N_{H}[a]\setminus K^{\prime}=N_{H^{\prime}}[u]$ and thus the set $(W\setminus\{a\})\cup\{u\}$ induces an obstruction in $H^{\prime}$, a contradiction. Similarly, if $|W\cap K^{\prime}|=\{a,b\}$ the set $(W\setminus\{a,b\})\cup\{u,v\}$ induces an obstruction in $H^{\prime}$, leading once again to a contradiction. Finally, since the biconnected components can be computed in $O(n+m)$ time (28), since block graphs can be recognized in $O(n+m)$ time and since true twins can be detected in $O(n+m)$ time (using modular decomposition (38)), both rules can be computed in linear time.
@@ 3.2 Reducing Block Graph branches
Conversely, let $F^{\prime}$ be a $k$-edition of $G^{\prime}$ into a block graph. Let $u$ be the attachment point of $B$ in $G$. Observe that since $B^{R}$ is connected and not empty, $u$ is not a cut-vertex of the block graph $B$ and its neighborhood in $B$ is either a maximal clique or a single vertex. Since $H^{\prime}=G^{\prime}\triangle F^{\prime}$ and $B$ are block graphs and $N_{B}(u)$ is a single vertex or a maximal clique of $B$, Lemma 2.6 ensures that $H=(H^{\prime},\{u\})\otimes(B^{R},N_{B}(u))$ is a block graph. We can observe that $H=G\triangle F^{\prime}$, hence $F^{\prime}$ is a $k$-edition of $G$ into a block graph. This concludes the proof.
@@
Figure 4: Illustration of the case where $p_{1}$ and $p_{2}$ are not in the same connected component of $H^{\prime}$ in the Lemma 3.21. The white circles correspond to critical cliques of $B$. The blue sets correspond to $N_{B}(P_{1})$ and $N_{B}(P_{2})$.
Let $F$ be a $k$-edition of $G$ into a block graph, $H=G\triangle F$ and $H^{\prime}=H\setminus V(B^{R})$. First consider the case in which $p_{1}$ and $p_{2}$ are in two different connected components in $H^{\prime}$. Let $H^{\prime}_{1},H^{\prime}_{2}$ be the two connected components of $H^{\prime}$ containing $p_{1}$ and $p_{2}$, respectively, and $H^{\prime}_{3}$ the remaining of $H^{\prime}$. It follows from Lemma 2.6 that the graph $H^{*}$ resulting from the disjoint union of
@@
We can assume that if $p_{1}$ and $p_{2}$ are in the same component of $H^{\prime}$, $F$ contains an edge-cut of $B$ or there would be a path $\pi_{B}$ from $p_{1}$ to $p_{2}$ in $B$ that still exists in $B\triangle F$. Let $M$ be a min-cut of $B$ and consider $B_{1}$ (resp. $B_{2}$) the connected component containing $p_{1}$ (resp. $p_{2}$) in $B\triangle M$. Both $B^{R}_{1}=B_{1}\setminus\{p_{1}\}$ and $B^{R}_{2}=B_{2}\setminus\{p_{2}\}$ are induced subgraphs (one of them possibly empty) of $B$ and block graphs by heredity. By Lemma 2.6, the following graph is a block graph:
@@
Figure 5: Illustration of the case where $p_{1}$ and $p_{2}$ are in the same connected component of $H^{\prime}$ and $F$ contains a min-cut of $B$ in Lemma 4.42. The white circles correspond to critical cliques of $B$. The blue sets correspond to $N_{B}(P_{1})$ and $N_{B}(P_{2})$. The red dotted edges between $B_{1}$ and $B_{2}$ are edges of an edge-cut of $B$ deleted by the edition.
@@
In all cases, there exists a $k$-edition $F^{*}$ of $G$ into a block graph such that the set pairs of $F^{*}$ containing a vertex of $V(B^{R})$ is either empty or exactly a min-cut of $B$. This concludes the proof.
Rule 3.4
@@
Let $F$ be a $k$-edition of $G$ that satisfies Lemma 3.21 and $H=G\triangle F$. If $F$ contains a min-cut of $B$, observe that $mc(B^{\prime})=mc(B)\leqslant k$ and consider $B_{i}^{\prime},i\in\{1,2\}$ the connected component of $B^{\prime}\triangle F_{m}$ containing $p_{i}$, where $F_{m}$ is a min-cut of $B^{\prime}$. By Lemma 2.6 the graph
@@
is a block graph. Let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$. Since $mc(B^{\prime})=mc(B)\leqslant k$, by construction $|F^{\prime}|=|F|$ , therefore $F^{\prime}$ is a $k$-edition of $G^{\prime}$. If $F$ does not contain a min-cut of $B$, then no vertex of $B^{R}$ is affected and $p_{1},p_{2}$ are not in the same connected component of $H\setminus V(B^{R})$. Let $H_{1}$ and $H_{2}$ be the connected components of $H\setminus V(B^{R})$ that contain respectively $p_{1}$ and $p_{2}$ and $H_{3}$ the remaining components of $H\setminus V(B^{R})$. By Lemma 2.6 the graph $H^{\prime}$ corresponding to the disjoint union of
@@
and $H_{3}$ is a block graph. We can observe that $H^{\prime}=G^{\prime}\triangle F$, hence $F$ is a $k$-edition of $G^{\prime}$ into a block graph. The other way is symmetrical.
@@ 3.3 Bounding the size of the kernel
Proof 3.26.
We show that we can enumerate all $1$-BG-branches and $2$-BG-branches in polynomial time. Recall that recognition of block graphs can be computed in linear time. All $1$-BG-branches can be enumerated by removing a vertex $v\in V(G)$ and looking among the remaining connected components those that induce a connected block graph together with $v$. We proceed similarly to detect clean $2$-BG-branches by removing a pair of non-adjacent vertices $u,v\in V(G)$ and looking among the remaining connected components those that induce a connected block graph together with $u$ and $v$.
@@
Proof 3.27.
Let $(G=(V,E),k)$ be a reduced yes-instance of Block Graph Editing, $F$ a $k$-edition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. Let $A$ be the set of affected vertices of $H$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T_{A}$ be the connected minimal induced subgraph of $H$ that spans $A$, and $d(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$. By Lemma 2.10, $|d(T_{A})\setminus A|\leqslant 3\cdot|A|\leqslant 6k$.
@@
Now consider the connected components of type 2. We can observe that such a component $K$ together with its adjacent vertices $x,y\in A\cup d(T_{A})$ form a clean 2-$BG$-branch, and thus contains at most $32k+$ 3 vertices. By Lemma 2.10, $T_{A}\setminus(A\cup d(T_{A}))$ contains at most $2\cdot|A|=4k$ connected components which is exactly the number of type 2 components, hence there are at most $4k\cdot(2k+1)$ vertices in connected components of type 2.
We conclude that if $(G,k)$ is a reduced yes-instance, then $|V(G)|\leqslant 24k^{2}+28k$. Therefore, given a reduced instance $(G,k)$ of Block Graph Editing, if $|V(G)|>24k^{2}+28k$ we can safely conclude that $(G,k)$ is a no-instance and return a trivial no-instance of constant size, thus proving the size of our kernel. Finally we claim that the reduced instance can be computed in polynomial time. Indeed, all reduction rules can be exhaustively applied in polynomial time (Lemmata 3.17 and 3.25).
@@
For the case of Block Graph Deletion, all the lemmata in this section are still true. In particular Lemma 3.21 holds by simply replacing ”$k$-edition” with ”$k$-deletion”, and the proof for the deletion variant is contained in the edition one. Using the same set of rules, we have a kernel with $O(k^{2})$ vertices for Block Graph Deletion.